I don't know how I got this wrong....One number is two more than a second number. Five times the first is 11 less than 4 times the second. Find the numbers.
I came up with the following equations:
x=2+y
5*X=4*Y-11
The value of the second number is 21. Apparently, this was wrong.
The value of the first number is 23.
Please explain.
Thanks.

Answer by mananth(16946)

x=2+y...............1
5x=4y-11.............2
Multiply (1) by 5
5x=10+5y

10+5y=4y-11
y=-21
x=2+y
x=2+(-21)
x=-19
-19,-21 are the numbers